# Faster Online Computation of String Net Frequency

### v1

@@ 3.5 Dealing with Type (ii) node
Let $c$ the first character of the edge from $av$ to $ax$. Then $(d,c)\in\Phi_{aS}(av)$ only if the node $ax$ is a leaf, where $ax$ is the destination node of the Weiner link from $v$ labeled $a$.
@@
Since $avc$ is a prefix of $ax$, $davc$ is a prefix of $dax$. Thus, due to Lemma 3, $(d,c)$ is the unique character pair for the character $c$ that can possibly be a member of $\Phi_{aS}(av)$.
If $ax$ is not a leaf, then $\mathsf{occ}_{aS}(avc)\geq\mathsf{occ}_{aS}(ax)>1$, implying $(d,c)\notin\Phi_{aS}(av)$. ∎
3.6 Dealing with Type (iii)-1 node
Here we show that the number of nodes in Type (iii)-1 is also at most one per iteration.
@@ 3.6 Dealing with Type (iii)-1 node
Let $v$ be the lowest ancestor of leaf $S$ in $\mathsf{STree}(S)$ such that ${\mathsf{W\_link}}_{a}({v})$ is defined. Then $v$ is the only possible node in Type (iii)-1.
@@
Suppose that $v$ is a node in Type (iii)-1, which implies that $\mathsf{occ}_{S}(av)=1$ and $\mathsf{occ}_{aS}(av)>1$. Let $ax$ be the destination node of the soft Weiner link ${\mathsf{W\_link}}_{a}({v})$. Then $av$ is a proper prefix of $ax$, and $av$ is an implicit node in $\mathsf{STree}(S)$ before the update. Since $\mathsf{occ}_{S}(av)=1$, we have $\mathsf{occ}_{S}(ax)=1$ which means that $ax$ is a leaf. Since the right-end of $S$ terminates with \$currency-dollar\\$\$, $x$ is also a leaf.
Consider the case where $x$ is a child of $v$. Then $(a,c)\in\Phi_{S}(v)$, where $c$ is the first character of the out-edge of $v$ leading to $x$ (see also the left diagram of Figure 2). Since $\mathsf{occ}_{aS}(av)>1$, $av$ becomes a branching node in $\mathsf{STree}(aS)$. Thus $(a,c)\notin\Phi_{aS}(v)$ holds for node $v$.
Let $v^{\prime}$ be any node in the path between $v$ and $z$ exclusive. Note that those nodes $v^{\prime}$ are only the other nodes than $v$ such that $\mathsf{occ}_{S}(av^{\prime})=1$ and $\mathsf{occ}_{aS}(v^{\prime})>1$. However, since their descendant $v$ is an internal node, $(a,c^{\prime})\notin\Phi_{aS}(v^{\prime})$ where $c^{\prime}$ is the only possible paring character for $a$ with node $v^{\prime}$.
Thus $v$ is the only possible node in Type (iii)-1. ∎
3.7 Dealing with Type (iii)-2 node
Here we show that the number of nodes in Type (iii)-2 is also at most one per iteration.
Lemma 10.
Let $z$ be the lowest ancestor of leaf $S$ in $\mathsf{STree}(S)$ such that ${\mathsf{W\_link}}_{a}({z})$ is hard. Then $az$ is the only possible node in Type (iii)-2.
Proof.
Suppose that $az$ is a node in Type (iii)-2, which implies that $\mathsf{occ}_{S}(aze)=1$ and $\mathsf{occ}_{aS}(aze)>1$ for a character $e$. Clearly node $az$ has an out-edge that begins with this character $e$. Since $\mathsf{occ}_{S}(aze)=1$, this edge leads to a leaf in $\mathsf{STree}(S)$. Moreover, this leaf is $ax$. Let $d$ be a unique character that immediately precedes $ax$ in $S$. Then, $(d,e)\in\Phi_{S}(az)$ (see also the left diagram of Figure 2).
Consider the case where $av$ becomes a branching node in $\mathsf{STree}(aS)$ after the update. Then, $av$ becomes a child of $az$, meaning that $\mathsf{occ}_{aS}(aze)>1$. Thus $(d,e)\notin\Phi_{aS}(az)$ for node $az$. Since Weiner’s algorithm inserts at most one new branching node, $az$ is the only node in Type (iii)-2. ∎
@@ 3.8 Complexities
It is known that Weiner’s algorithm finds the nodes $u$, $v$, and $z$ in $O(1)$ amortized time for a new character $a$ that is prepended to $S$ (Lemma 5). It follows from our previous discussions that given the nodes $u$, $v$, and $z$, we can compute $\mathsf{NF^{+}}(aS)\triangle\mathsf{NF^{+}}(S)$ in $O(\log\sigma)$ time, where the $\log\sigma$ factor is the cost to access the Weiner links.
@@
By maintaining a list for $\mathsf{NF^{+}}(S)$ in an online manner, we obtain the following:
@@ Acknowledgements
This work is supported by JSPS KAKENHI Grant Numbers JP20H05964, JP23K24808, JP23K18466.
The author thanks Takuya Mieno for fruitful discussions.

### v2

@@ 3.6 Dealing with Type (iii)-1 node
Let $v$ be the lowest ancestor of leaf $S$ in $\mathsf{STree}(S)$ such that ${\mathsf{W\_link}}_{a}({v})$ is defined (see also Figure 3), and let $ax$ be the destination node of ${\mathsf{W\_link}}_{a}({v})$ from $v$ labeled $a\in\Sigma$ in $\mathsf{STree}(S)$ before the update. Let $c$ be the first character of the edge from $v$ to $x$. Then $v$ is the only possible node of Type (iii)-1. Also, $(a,c)$ is the only possible character pair such that $(a,c)\in\Phi_{S}(v)\setminus\Phi_{aS}(v)$.
@@
Consider the case that $\mathsf{occ}_{S}(av)=1$ and $\mathsf{occ}_{aS}(av)>1$. Since $\mathsf{occ}_{S}(av)=1$, we have $\mathsf{occ}_{S}(ax)=1$ and thus $ax$ is a leaf. Since $ax\in\mathsf{Suffix}(S)$, $x\in\mathsf{Suffix}(S)$ and hence $x$ terminates with \$currency-dollar\\$\$. Since $v$ is an internal node, $v$ cannot end with \$currency-dollar\\$\$ and thus $v$ is a proper prefix of $x$. Thus $av$ is on the edge leading to the leaf $ax$ in $\mathsf{STree}(S)$ before the update. Since $x$ terminates with \$currency-dollar\\$\$, $x$ is also a leaf. Let $c$ be the first character of the path label from $v$ toward $x$. If $x$ is not a child of $v$, then $\mathsf{occ}_{S}(vc)>1$ and thus $v$ cannot be of Type (iii)-1. Consider the case where $x$ is a child of $v$. In this case, $\mathsf{occ}_{S}(av)=\mathsf{occ}_{S}(vc)=\mathsf{occ}_{S}(avc)=1$ (see also the left diagram of Figure 3), and thus we have $(a,c)\in\Phi_{S}(v)$. After the update, $av$ becomes a branching node in $\mathsf{STree}(aS)$, meaning that $\mathsf{occ}_{aS}(av)>1$ (see also the right diagram of Figure 3). By Lemma 6, $(a,c)\notin\Phi_{aS}(v)$ holds for node $v$. This can also be observed with the corresponding existing substructure in form of Figure 2 that disappears in $\mathsf{STree}(aS)$. Overall, if the aforementioned conditions are satisfied, then $v$ is a node of Type (iii)-1. Since $\mathsf{occ}_{S}(avc)=1$, $c$ is the only right-extension of $av$ in $S$. It now follows from Lemma 3 that $a$ is the only pairing character for $c$, and hence $(a,c)$ is the only possible character pair such that $(a,c)\in\Phi_{S}(v)\setminus\Phi_{aS}(v)$.
Next, we show that $v$ is the only possible node that is of Type (iii)-1. Let $v^{\prime}$ be any node in the path between $v$ and $z$ exclusive, where $z$ is the lowest ancestor of the leaf $S$ that has a hard Weiner link ${\mathsf{W\_link}}_{a}({z})$ labeled $a$ in $\mathsf{STree}(S)$ before the update. Note that those nodes $v^{\prime}$ are the only nodes other than $v$ such that $\mathsf{occ}_{S}(av^{\prime})=1$ and $\mathsf{occ}_{S}(v^{\prime})>1$. However, since their descendant $v$ is an internal node, we have $\mathsf{occ}_{S}(v^{\prime}c^{\prime})>1$ and thus $(a,c^{\prime})\notin\Phi_{aS}(v^{\prime})$ holds, where $c^{\prime}$ is the only possible paring character for $a$ with node $v^{\prime}$ due to Lemma 3. Thus $v$ is the only possible node of Type (iii)-1. ∎
The node $v$ is found by climbing up the path from the leaf for $S$, during the update from $\mathsf{STree}(S)$ to $\mathsf{STree}(aS)$. The node $ax$ is obtained by the Weiner link ${\mathsf{W\_link}}_{a}({v})$ from $v$, and the node $x$ is obtained by the reversed Weiner link (i.e. the suffix link) from $ax$. The character $c$ is obtained by using the string depths of the nodes $av$ and $ax$.
Recall that $z$ is the lowest ancestor of the leaf $S$ that has a hard Weiner link ${\mathsf{W\_link}}_{a}({z})$ labeled $a$ in $\mathsf{STree}(S)$ before the update. If $v\neq z$, then $av$ is on an edge. We can easily check whether $x$ and $ax$ are leaves and $x$ is a child of $v$ in $\mathsf{STree}(S)$. Notice that the first condition $v\neq z$ is critical, since if $v=z$, then $av$ ($=az$) is an internal node in $\mathsf{STree}(S)$ and thus $\mathsf{occ}_{S}(av)>1$. Now we have two sub-cases:
  • If $\phi_{S}(v)=1$ (namely $\Phi_{S}(v)=\{(a,c)\}$), then we get $\phi_{aS}(v)=0$ (namely $\Phi_{aS}(v)=\emptyset$) after the update, and thus $(u,\phi_{S}(v))$ is removed from the list $\mathcal{L}$.
  • If $\phi_{S}(v)>1$, then we update $\Phi_{aS}(v)\leftarrow\Phi_{S}(v)\setminus\{(a,c)\}$ and replace $(v,\phi_{S}(v))$ with $(v,\phi_{S}(v)-1)=(v,\phi_{aS}(v))$ in $\mathcal{L}$.
3.7 Dealing with Type (iii)-2 node
Here we show that the number of nodes of Type (iii)-2 is also at most one per iteration.
@@ 3.7 Dealing with Type (iii)-2 node
Let $z$ be the lowest ancestor of leaf $S$ in $\mathsf{STree}(S)$ such that ${\mathsf{W\_link}}_{a}({z})$ is hard (see also Figure 3). Let $v$ be the lowest ancestor of leaf $S$ in $\mathsf{STree}(S)$ such that ${\mathsf{W\_link}}_{a}({v})$ is defined, and let $ax$ be the destination node of ${\mathsf{W\_link}}_{a}({v})$ from $v$ labeled $a\in\Sigma$ in $\mathsf{STree}(S)$ before the update. Let $e$ be the first character in the path from $az$ to $ax$. Then $az$ is the only possible node of Type (iii)-2. Also, $(d,e)$ is the only possible character pair such that $(d,e)\in\Phi_{S}(az)\setminus\Phi_{aS}(az)$, where $d\in\Sigma$ is the unique character which immediately precedes $ax$ in $S$.
@@
Recall that $az$ is the parent of $ax$ in $\mathsf{STree}(S)$ before the update. Consider the case that $\mathsf{occ}_{S}(aze)=1$, where $ax$ is a leaf in $\mathsf{STree}(S)$. Also, consider the case that $\mathsf{occ}_{S}(daz)=1$. Since $daze$ is a prefix of $dax$, $\mathsf{occ}_{S}(daze)\geq 1$. Moreover, since $\mathsf{occ}_{S}(aze)=1$, we have $\mathsf{occ}_{S}(daze)=1$. Therefore, $(d,e)\in\Phi_{S}(az)$ by Lemma 6. Consider the case where $av$ becomes a branching node in $\mathsf{STree}(aS)$ after the update. Then, $av$ becomes a child of $az$, meaning that $\mathsf{occ}_{aS}(aze)>1$. Thus $(d,e)\notin\Phi_{aS}(az)$ holds for $az$. Overall, if the aforementioned conditions are satisfied, then $az$ is a node of Type (iii)-2. This can also be observed with the corresponding existing substructure in form of Figure 2 that disappears in $\mathsf{STree}(aS)$.
Since Weiner’s algorithm inserts at most one new internal node, $az\in\mathsf{M}(S)$ and $e\in\Sigma$ are the only maximal repeat and character that can satisfy $\mathsf{occ}_{S}(aze)=1$ and $\mathsf{occ}_{aS}(aze)>1$. By Lemma 3, $(d,e)$ is the the only possible character pair such that $(d,e)\in\Phi_{S}(az)\setminus\Phi_{aS}(az)$. ∎
The node $az$ is found via the hard Weiner link ${\mathsf{W\_link}}_{a}({z})$ in $\mathsf{STree}(S)$ from $z$, during the update from $\mathsf{STree}(S)$ to $\mathsf{STree}(aS)$. The character $e$ is found by using the string depths of the nodes $az$ and $ax$. We can find the character $d$ and can check whether $ax$ is a leaf or not analogously to the case of Type (iii)-1 nodes. The rest is also analogous to Type (iii)-1 nodes.
@@ 3.8 Complexities
It is known that Weiner’s algorithm finds the nodes $u$, $v$, and $z$ in $O(1)$ amortized time for a new character $a$ that is prepended to $S$ (Lemma 5). It follows from our previous discussions that given the nodes $u$, $v$, and $z$, we can compute $\mathsf{NF^{+}}(aS)\triangle\mathsf{NF^{+}}(S)$ in $O(\log\sigma)$ amortized time, where the $\log\sigma$ factor is the cost to access the Weiner links from nodes $v$ and $z$.
@@
By maintaining the list $\mathcal{L}$ for $\mathsf{NF^{+}}(S)$ in an online manner, we obtain the following:
@@ Acknowledgements
This work is supported by JSPS KAKENHI Grant Numbers JP23K24808 and JP23K18466. The author thanks Takuya Mieno for fruitful discussions.